B Excel #1

Proposal #1 would extend trade credit to some customers that previously have been denied credit because they were considered poor risks.   Sales are projected to increase by $120,000 per year if credit is extended to these new customers. Of the new accounts receivable generated, 6% are projected to be uncollectible. Additional collection costs are projected to be 5% of incremental sales, and production and selling costs are projected to be 80% of sales. Your firm expects to pay a total of 30% of its income after expenses in taxes. (Show Work)

  1. Compute the incremental income after taxes that would result from these projections:
  1. Compute the incremental Return on Sales if these new credit customers are accepted:

If the receivable turnover ratio is expected to be 4 to 1 and no other asset buildup is needed to serve the new customers…

  1. Compute the additional investment in Accounts Receivable
  2. Compute the incremental Return on New Investment
  1. If your company requires a 20% Rate of Return on Investment for all proposals, do the numbers suggest that trade credit should be extended to these new customers? Explain.
Answer #1

Given:

incremental Sales = $120,000
uncollectibles sales = 6%* sales
Collection cost = 5%*sales
Production and selling cost = 80%*sales
Income tax = 30%

a) Incremental income after tax is calculated as below:

Incremental Sales 120000
Less Uncollectible sales of 6% 7200
Less collection cost 6000
Less Production and selling cost 96000
Earning before tax 10800
Less tax @ 30% 3240
Net Incremental Income after tax 7560


Incremental Income tax $7560

b) Incremental return on sales = Net Income / Sales
                                                 = 7560 / 120000

                                                 = 6.30%
c)Receivable turnover ratio = 4:1
Accounts receivables = Sales / Receivable turnover ratio
                                   = 120000 / 4
                                   = $30000

Additional investment in accounts receivables is $30000

c) Incremental return on investment

Only incremental investment is accounts receivables. Therefore
Return on investment = Net Income / Accounts receivables
                                      = 7560 / 30000
                                     
                                      = 25.20%

d) Yes, trade credit could be given to them as the return on investment 25.20% is higher than the 20% required rate of return